A bag contains 9 red marbles, 8 white marbles, and 6 blue marbles. Randomly choose two marbles, one at a time, and without replacement. Find the following:
a) The probability that the first marble is red and the second is white.
b) The probability that both are the same color.
c) The probability that the second marble is blue.

Found 2 solutions by ewatrrr, Boreal:
Answer by ewatrrr(24785)

23 marbles: 9 red marbles, 8 white marbles, and 6 blue marbles
Draw two, w/o replacement
a) The probability that the first marble is red and the second is white.
P = (9/23)(8/22)
b) The probability that both are the same color.
P = 2R + 2W + 2B
P = (9/23)(8/22) + (8/23)(7/22) + (6/23)(5/22)
c) The probability that the second marble is blue.
P = (17/23)(6/22) + 2B (Need to consider both are Blue)
P = .202 + (6/23)(5/22) = .202 + .059 = .261

Answer by Boreal(15235)

23 marbles
9/23*8/22, since there is one less marble
That is 72/506 or 0.142
-------------------------------
same color
9/23*8/22 (red)
8/23*7/22 (white)
6/23*5/22 (blue)
The denominator is 506
The numerator is 72+56+30=158
158/506=0.312
----------------------------------------
Second marble blue:
3 possibilities
first red and then it is 9/23*6/22
first blue and then it is 6/23*5/22
first is white and then it is 8/23*6/22
Denominator is 506
Numerator is 54+30+48=132
132/506=0.261